OpenStudy (anonymous):

I needd helppp? Identify the correct slope and y-intercept of the equation 5x - 3y = -6 a. slope = -5; y-intercept = (0, -6) b. slope = 5/3; y-intercept = (0, 2) c. slope = -3/5; y-intercept = (0, 2) d. slope = 5/3; y-intercept = (0, -6)

OpenStudy (blurbendy):

B) 5x - 3y = -6 -3y = -5x -6 y = 3/5x +2

OpenStudy (anonymous):

shannkkkuuuuu

OpenStudy (blurbendy):

np!